Could someone show me how to solve this word problem for triangles?
The base of the triangle is 7cm greater then the height. The area is 39cm^2. Find the height and the length of the base.
: Could someone show me how to solve this word problem for triangles?
The base of the triangle is 7cm greater then the height. The area is 39cm^2. Find the height and the length of the base.

Answer by oscargut(667)

b=h+7
bh/2=39 yhen bh=78
then
(h+7)h=78
h^2+7h-78=0


h = (-7 +- sqrt( 7^2-4*1*78 ))/2 =
(-7 +- sqrt(361))/2 =(-7 +- 19)/2
then h=(-7+19)/2=6
Answer:
h=6 and b=13